# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v2

@@ II.3 CSS Floquet code
Having discussed the honeycomb codes, we now give a brief review of another type of Floquet code, the CSS Floquet code, which was first presented in Refs. [18, 13]. We refer the readers to Appendix A.2 for more details on the code. The CSS Floquet code is defined on the same honeycomb lattice. We show the code’s measurement cycle in Fig. 1(a). We measure operators defined on edges using the $r\rightarrow g\rightarrow b$ cycle, but alternate between measuring $XX$ and $ZZ$ operators on these edges. Hence, this code has a period-6 measurement cycle, and its logical operator evolutions also have period-6 (see Appendix A.2 for an illustration). Even though a honeycomb code’s measurement schedule has only a period of 3, for consistency, we will define 1 QEC round to be 6 measurement subrounds for all codes studied in this paper. The check operators, stabilisers and logicals all are either $X$- or $Z$-type and, for this reason, the code is called the CSS Floquet code.
Unlike the honeycomb codes, the CSS Floquet code has no persistent stabiliser operators. Instead, each plaquette stabiliser is repeatedly reinitialised and measured out, with subsequent measurement values compared to form a detector. In contrast to the honeycomb codes where the values of the plaquette stabilisers are inferred from measurements of check operators in two consecutive subrounds, here the plaquette operator measurements take only a single subround [13, 18]. Because of these single-step measurements, at each measurement subround, there is always one type of plaquette operator that anti-commutes with the check measurements, and hence their values are undetermined (see Appendix A.2). These “inactive” plaquettes, however, will be reinitialized at the next measurement subround. As a result, at each time step, the ISG contains $X$-type and $Z$-type stabiliser operators defined on two of the three colours of plaquette and either an $X$-type or a $Z$-type operator defined on the other colour. For instance, after measuring the red-$X$ checks, the ISG contains $X^{\otimes 6}$ and $Z^{\otimes 6}$ blue and green plaquette operators, but only $X^{\otimes 6}$ red plaquette operators, since $Z^{\otimes 6}$ red plaquette operators anti-commute with the red-$X$ checks (see Appendix A.2 for more details).
@@
Figure 2: Persistent symmetry of the X3Z3 Floquet code leads to improved performance under $Z$-biased noise. (a) An example of such a symmetry along a single unshaded strip of the lattice, together with a Pauli-$Z$ error, is shown throughout the measurement cycle. Arrows indicate the type of check ($c\text{A}$ or $c$B for some colour $c$) measured at each subround. A-type plaquette operators whose product forms the symmetry at that subround are highlighted in darker colours. (b) An example of a symmetry after $r$A checks are measured is the product of red and blue A-type plaquettes shown. Since this product is the identity along the vertical unshaded strip and commutes with $Z$ errors on all qubits, the syndromes/anyons [represented by coloured circles in (a)] must appear in pairs along the vertical strips.
Figure 3: The A-type detector graph of the X3Z3 Floquet code under pure dephasing noise. Only a single vertical strip [e.g., that supporting the symmetry in Fig. 2(b)] is shown. The illustrated part of the graph is disconnected from that of any neighbouring strips. Arrows represent only the A-type measurements. The B-type subrounds are not shown, since they do not influence the detectors in this graph. Each node (representing a detector) is placed in the center of a plaquette and at the subround in which the corresponding detector is formed. For example, the bottom layer corresponds to the blue A-check measurements, at which time the red A-type plaquettes are measured.
III X3Z3 Bias-tailored Floquet Code
@@
The decoding of the X3Z3 Floquet code under $Z$-biased noise is simplified by the presence of a symmetry in the decoding graphs, which we will explain below. This is a space-time analogue of the symmetries present in bias-tailored static codes such as the XZZX code [41] and the domain wall colour code [45]. In such codes, under the infinitely-biased code-capacity noise model (with only single-qubit $Z$ errors, for example), syndromes are forced to come in pairs along one-dimensional (1D) strips of the lattice. This results from strips of stabilisers multiplying together to the identity along one domain and commuting with all $Z$ errors on all qubits: these are 1D symmetries of the stabiliser code under infinitely biased noise. We will refer to (plaquette) stabilisers in the ISG that are flipped by an error as anyons. In bias-tailored static codes, anyons can propagate within strips but cannot move outside the strip without changing their Pauli type. We will show that there also exists a similar symmetry in the bias-tailored Floquet code.
@@
III.1 Persistent symmetry and two-dimensional decoding under biased noise
@@
There is no constant stabiliser group when viewing the X3Z3 Floquet code as a subsystem code (in a subsystem code, the stabiliser group is defined as the center of the gauge group) [13, 18]. Surprisingly, despite this fact, at every time step there do exist operators that form a symmetry under pure dephasing noise (without measurement errors). An example of such a symmetry is shown in Fig. 2 which is a symmetry on the unshaded strips formed by the A-type plaquettes. A similar symmetry can also be found on the shaded strips which are formed by B-type plaquettes. These symmetries, which are formed by plaquette operators of a single type along vertical strips of the lattice at time step $t$, also commute with all operators in $\mathcal{M}_{t+1}$ (the set of check measurements to be performed in the next subround). To see why the symmetries are formed at each measurement subround, consider one particular time step as shown in Fig. 2(b), i.e., after the red A-type measurements. As depicted, the product of red and blue A-type plaquettes is the identity on all qubits in the unshaded vertical strip, and yields Pauli $Z$ operators on some qubits in the neighbouring shaded vertical strips. At this subround, there is a similar symmetry that has the identity on the shaded strips, which is formed by the product of blue and green B-type plaquettes.
@@
As a result of each of these symmetries, Pauli-$Z$ errors create anyons in pairs along vertical strips. An example of such anyons is shown in each measurement subround of Fig. 2(a). Note that the anyons change plaquette locations between subrounds, but always respect the symmetry. The anyons depicted in Fig. 2 represent detectors that would be triggered if the Pauli-$Z$ error shown occurred at that time step. For example, if the $Z$ error occurs immediately after the red A-check measurements, it will trigger a red A detector after subsequent blue A-check measurements and a blue A detector after green A-check measurements.
@@
Unlike the symmetry in static stabiliser codes, this symmetry does not allow for 1D decoding even when measurement errors are absent. While in static codes, only measurement errors produce “time-like” edges, in Floquet codes, even Pauli errors produce time-like edges (between detectors formed at different times). To demonstrate this, we display in Fig. 3 a portion of the A-type detector graph under infinitely-biased single-qubit noise with no measurement errors. As can be seen, there is a disconnected subgraph defined along one strip of the code lattice. Even considering this simple noise model, the decoding graph in the infinite bias regime is two-dimensional (i.e., the graph is planar). This results from the fact that neighbouring plaquettes are measured at different times. We emphasise that although the above discussion is based on $Z$ errors, the same analysis also holds for $X$ errors because $X$ and $Z$ operators are interchangeable for the X3Z3 code. As a result, the performance of the X3Z3 code under $X$-biased noise is expected to be the same as that under the $Z$-biased noise model.
III.2 Bias-preserving parity-check circuits
Figure 4: Depth-2 check measurement circuits using CNOT gates along with resets and single-qubit measurements in the $Z$, $X$ and $Y$ bases for (a) $Z_{1}Z_{2}$, (b) $X_{1}X_{2}$, and (c) $Z_{1}X_{2}$ checks, respectively.
@@ III.2 Bias-preserving parity-check circuits
While two-body measurements are native to certain architectures, for example, Majorana [10, 56, 57] and photonic [58, 17] platforms, most architectures require quantum circuits to carry out such measurements. To maintain the high-performance of bias-tailored codes in these hardwares, the parity-check measurement circuits need to be constructed in a bias-preserving manner, such that they do not propagate frequently-occurring errors to rarely-occurring errors. To this end, we design two-qubit parity check circuits that preserve the $Z$ bias on data qubits. That is, the probability of $X$ and $Y$ errors on data qubits after these circuits remains small (proportional to single-qubit $X$ or $Y$ error probabilities). Our bias-preserving parity check measurement circuits areconstructed by using an ancilla circuit connecting two data qubits, which can be realised even in devices with minimal connectivity, such as the heavy-hexagonal layout [53, 52].
@@
For a circuit to be bias preserving, it can be constructed using only gates that do not change the error type under conjugation. For example, one way to construct bias-preserving circuits is to use only CNOT gates in the measurement circuits as they only propagate errors to others with the same Pauli type. For the X3Z3 code, we need to construct bias-preserving circuits for three different kinds of parity checks: $ZZ$, $XX$ and $ZX$. All these three circuits can be constructed using only CNOT gates as shown in Fig. 4. As depicted, these circuits also include resets and measurements of the ancilla qubit in the Pauli bases, i.e., $Z$, $X$ and $Y$ bases for $ZZ$, $XX$ and $ZX$ checks, respectively. Note that the circuits are still bias-preserving even if the resets/measurements in the $X$ and $Y$ bases are compiled in terms of resets/measurements in the $Z$ basis with additional single-qubit Clifford gates. This is because a $Z$ error on any qubit at every time step in the compiled circuits still propagates to data qubits as a $Z$ error, or as a check operator about to be measured.
@@
The parity-check circuits shown in Fig. 4, however, may not preserve the noise bias if errors happen during the CNOT gates. This is because a $Z$ error occurring on the target qubit during the application of a conventional (non-bias preserving) CNOT gate may result in a combination of $Z$ and $X$ errors on the target qubit after the CNOT gate [59]. One way to remedy this problem is to use bias-preserving CNOT gates [59]. These gates, however, are available only in specific platforms such as cat qubits. Here, we propose a more general approach to constructing $Z$-bias preserving parity check circuits without requiring the bias-preserving gates. While our proposed circuits can preserve only a specific type of noise bias, i.e., the $Z$ bias, these circuits are built using conventional (which may not be bias-preserving) two-qubit gates and thus can be implemented in many different architectures.
Our proposed fully bias-preserving parity-check circuits are depicted in Fig. 5. These circuits are constructed using either CZ gates or CNOT gates targeting ancilla qubits, which only propagate Pauli errors on the ancilla qubits as $Z$ errors on the data qubits. While mid-gate $Z$ errors may result in $X$ errors on the ancilla qubits, such errors do not degrade the bias of errors affecting the data qubits. We note that while the $ZZ$ parity check measurements are performed by reading out the ancilla qubits like in standard syndrome extraction circuits, the $XX$ and $ZX$ parity values are obtained by measuring the data qubits after some two-qubit gates are applied on them. These two qubits gates are required for the parity measurements so as not to reveal each individual data qubit’s state. For example, the two mid-circuit $X$-measurements in Fig. 5(b), are equivalent to reading out the $X_{1}X_{a}$ and $X_{2}X_{a}$ parity values, where $X_{a}$ acts on the ancilla qubit which is prepared in the |$\ket{0}$ state. The final $Z_{a}$ measurement disentangles the ancilla qubit, leaving the data qubits in an eigenstate of $X_{1}X_{2}$. The circuit therefore performs the required $XX$ check measurement.
The mixed-type $Z_{1}X_{2}$ parity circuit can be constructed using CZ gates between the data qubits, as shown in Fig. 5(c). If the connections between the data qubits are not available in the device such as in hardwares with a heavy-hexagonal lattice, these CZ gates have to be implemented using ancilla qubits in-between the data qubits. We show such an implementation of next-nearest-neighbour CZ gates in the right-hand side of Fig. 5(c) (this is adapted from Ref. [52]). In the figure, we also present optional measurements, coloured in grey, which provide “flag information” for detecting $Z$ errors on the measurement ancilla that may have propagated to the data qubits.
We emphasise that, even if the $X$ measurements in the parity-check circuits above are implemented using $Z$ measurements sandwiched by Hadamard gates, the circuits are still bias preserving since a $Z$ error occurring between those Hadamard gates is harmless. This is because a $Z$ error happening before the measurement is immediately absorbed by the $Z$ measurement without flipping its outcome, and a $Z$ error after the measurement is a stabiliser of the state (up to a sign), so does nothing. Moreover, the circuit will not change the noise bias significantly even if the Hadamard gates are noisy. The reason is because in many architectures, the single-qubit gate errors are not the predominant error and are usually much smaller than the two-qubit gate errors [60, 61, 62, 63].
@@
Figure 5: Measurement circuits for (a) $ZZ$, (b) $XX$, and (c) $Z_{1}X_{2}$ checks that preserve the $Z$ noise bias on data qubits. Data qubits are labelled by Data1 and Data2, while the middle qubit in each circuit is a measurement ancilla. In (c), the equality can be checked by noting that the depth-3 circuit before and after the mid-circuit measurement performs a next-nearest-neighbour CZ gate, if the measurement ancilla is initially in the $\ket{+}$ state. The grey measurements can be optionally included to provide flag information.
IV Methods

### v3

@@ II.3 CSS Floquet code
Having discussed the honeycomb codes, we now give a brief review of another type of Floquet code, the CSS Floquet code, which was first presented in Refs. [18, 13]. We refer the readers to Supplementary Sec. IB of the Supplementary Information [61] for more details on the code. The CSS Floquet code is defined on the same honeycomb lattice. We show the code’s measurement cycle in Fig. 1(a). We measure operators defined on edges using the $r\rightarrow g\rightarrow b$ cycle, but alternate between measuring $XX$ and $ZZ$ operators on these edges. Hence, this code has a period-6 measurement cycle, and its logical operator evolutions also have period-6 (see Supplementary Fig. 2 of the Supplementary Information [61] for an illustration). Even though the honeycomb code’s measurement schedule has only a period of 3, for consistency, we will define 1 QEC round to be 6 measurement subrounds for all codes studied in this paper. The check operators, stabilisers and logicals all are either $X$- or $Z$-type and, for this reason, the code is called the CSS Floquet code.
Unlike the honeycomb codes, the CSS Floquet code has no persistent stabiliser operators. Instead, each plaquette stabiliser is repeatedly reinitialised and measured out, with subsequent measurement values compared to form a detector. In contrast to the honeycomb codes where the values of the plaquette stabilisers are inferred from measurements of check operators in two consecutive subrounds, here the plaquette operator measurements take only a single subround [13, 18]. Because of these single-step measurements, at each measurement subround, there is always one type of plaquette operator that anti-commutes with the check measurements, and hence their values are undetermined (see Supplementary Sec. IB of the Supplementary Information [61]). These “inactive” plaquettes, however, will be reinitialised at the next measurement subround. As a result, at each time step, the ISG contains $X$-type and $Z$-type stabiliser operators defined on two of the three colours of plaquette and either an $X$-type or a $Z$-type operator defined on the other colour. For instance, after measuring the red-$X$ checks, the ISG contains $X^{\otimes 6}$ and $Z^{\otimes 6}$ operators on both blue and green plaquettes, but only $X^{\otimes 6}$ red plaquette operators, since $Z^{\otimes 6}$ red plaquette operators anti-commute with the red-$X$ checks (see Supplementary Sec. IB of the Supplementary Information [61] for more details).
@@
II.4 X3Z3 bias-tailored Floquet code
@@
Figure 2: Persistent stabiliser-product symmetry of the X3Z3 Floquet code leads to improved performance under $Z$-biased noise. (a) An example of such a symmetry along a single unshaded strip of the lattice, together with a Pauli-$Z$ error, is shown throughout the measurement cycle. Arrows indicate the type of check ($c\text{A}$ or $c$B for some colour $c$) measured at each subround. A-type plaquette operators whose product forms the symmetry at that subround are highlighted in darker colours. Plaquettes indicated by coloured dots are the locations of syndromes that would be triggered if the $Z$ error shown were to occur at that measurement subround. (b) An example of a stabiliser-product symmetry just before $g$B checks are measured is the product of red and blue A-type plaquettes shown. Since this product is the identity along the vertical unshaded strip and commutes with $Z$ errors on all qubits, the syndromes/anyons [coloured dots in (a)] must appear in pairs along the vertical strips.
Figure 3: The 2D A-type detector graph of the X3Z3 Floquet code under pure dephasing noise. Only a single vertical strip [e.g., that supporting the stabiliser-product symmetry in Fig. 2(b)] is shown. The illustrated part of the graph is disconnected from those of neighbouring strips. Arrows represent only the A-type measurements. The B-type subrounds are not shown, since they do not influence the detectors in this graph. Each node (representing a detector) is placed in the centre of a plaquette and at the subround in which the corresponding detector is formed. For example, the bottom layer corresponds to the blue A-check measurements, at which time the red A-type plaquettes are measured.
The decoding of the X3Z3 Floquet code under $Z$-biased noise is simplified by the presence of a symmetry in the decoding graphs, which we will explain below. This is a space-time analogue of the symmetries present in bias-tailored static codes such as the XZZX code [42] and the domain wall colour code [46]. In such codes, under the infinitely phase-biased code-capacity noise model (with only single-qubit $Z$ errors, for example), syndromes are forced to come in pairs along one-dimensional (1D) strips of the lattice. This results from strips of stabilisers multiplying together to the identity along one domain and commuting with all $Z$ errors on all qubits: these are 1D symmetries of the stabiliser code under infinitely phase-biased noise. We will refer to (plaquette) stabilisers in the ISG that are flipped by an error as anyons. In bias-tailored static codes, anyons can propagate within strips but cannot move outside the strip without changing their Pauli type. We will show that there also exists a similar symmetry in the bias-tailored Floquet code.
@@
II.5 Persistent stabiliser-product symmetry and two-dimensional decoding of the X3Z3 Floquet code under biased noise
@@
There is no constant stabiliser group when viewing the X3Z3 Floquet code as a subsystem code (in a subsystem code, the stabiliser group is defined as the centre of the gauge group, which, in our case, is generated by all edge operators) [13, 18]. Surprisingly, despite this fact, at every time step there do exist operators that form a symmetry under pure dephasing noise (without measurement errors). An example of such a symmetry is shown in Fig. 2. It is a symmetry on one of the unshaded domains and it is formed by the product of A-type plaquettes. A similar symmetry can also be found on the shaded strips which are formed by the product of B-type plaquettes.
[…]
$\prod_{S\in\text{ISG}_{(t,t+1)}^{A,\mathcal{D}}}S=\left(\bigotimes_{k\in\mathcal{D}}I_{k}\right)\left(\bigotimes_{j\in\bar{\mathcal{D}}_{Z,t}}Z_{j}\bigotimes_{j^{\prime}\in\bar{\mathcal{D}}_{I,t}}I_{j^{\prime}}\right),$ (2)
where $\bar{\mathcal{D}}_{Z,t}$ and $\bar{\mathcal{D}}_{I,t}$ are some subsets of qubits in the domains adjacent to domain $\mathcal{D}$ on which the product of the A-type plaquette stabilisers yields Pauli $Z$ and the identity, respectively [for the example given in Fig. 2(b), the domain $\cal{D}$ is the unshaded vertical strip, the domains $\bar{\mathcal{D}}_{Z,t}$ and $\bar{\mathcal{D}}_{I,t}$ comprise the qubits in the neighbouring shaded vertical strips with $Z$-circled label and no label, respectively]. We use a subscript $t$ in the notations $\bar{\mathcal{D}}_{Z,t}$ and $\bar{\mathcal{D}}_{I,t}$ to indicate that the qubits belonging to these two subsets depend on the time step $t$. Note that we suppress the identity factors outside of $\mathcal{D}$ and its adjacent domains. In particular, by restricting this operator to domain $\mathcal{D}$, we have the following persistent stabiliser-product symmetry:
$\prod_{S\in\text{ISG}_{(t,t+1)}^{A,\mathcal{D}}}S|_{\mathcal{D}}$ $=\bigotimes_{k\in\mathcal{D}}I_{k},$ (3)
which implies that the error syndrome obeys a conservation law in the domain $\mathcal{D}$: the syndrome comes in pairs along this strip. At each subround, there also exists a similar symmetry on the shaded domains, which is formed by the product of B-type plaquettes. Therefore, for the pure $Z$ Pauli noise model, we can perform minimum-weight perfect matching decoding within each domain.
@@
To explain this more concretely, we use the anyon picture. These anyons, shown in Fig. 2(a), are interpreted as having locations given by the detectors that would be flipped if the corresponding error were to occur at that time step. For example, if the $Z$ error shown in Fig. 2(a) occurs just before the green B-check measurements (top-left panel of the figure), it will trigger a red A detector after two measurement subrounds and a blue A detector after a further two subrounds. As a result of each of the symmetries, Pauli-$Z$ errors create anyons in pairs along each vertical strip. An example of such anyons, which are formed due to a Pauli $Z$ error at a particular measurement subround, is shown in Fig. 2(a). While the plaquette locations of the anyons, triggered by the $Z$ error occurring on the same qubit, change every other subround, they always respect the symmetry, i.e., are aligned along the domain (see Fig. 2).
@@
Unlike the symmetry in static stabiliser codes, this stabiliser-product symmetry does not allow for 1D decoding even when measurement errors are absent. While in static codes, only measurement errors produce “time-like” edges, in Floquet codes, even Pauli errors produce time-like edges (between detectors formed at different times). To demonstrate this, we display in Fig. 3 a portion of the A-type detector graph under infinitely phase-biased single-qubit noise with no measurement errors. As can be seen, there is a disconnected subgraph defined along one unshaded domain of the code lattice. Even considering this simple noise model, the decoding graph in the infinite bias regime is two-dimensional (i.e., the graph is planar). This results from the fact that neighbouring plaquettes are measured at different times. We emphasise that although the above discussion is based on $Z$ errors, the same analysis also holds for $X$ errors because $X$ and $Z$ operators are interchangeable for the X3Z3 code. As a result, the performance of the X3Z3 code under $X$-biased noise is expected to be the same as that under the $Z$-biased noise model.
II.6 Bias-preserving parity-check circuits
@@
While two-body measurements are native to certain architectures, for example, Majorana [10, 62, 63] and photonic [64, 17] platforms, most architectures require quantum circuits to carry out such measurements. To maintain the high-performance of bias-tailored codes in these hardware, the parity-check measurement circuits need to be constructed in a phase-bias preserving manner, such that they do not propagate frequently-occurring errors to rarely-occurring errors. To this end, we design two-qubit parity check circuits that preserve the $Z$ bias on data qubits. That is, the probability of $X$ and $Y$ errors on data qubits after these circuits remains small (proportional to single-qubit $X$ or $Y$ error probabilities). Our bias-preserving parity check measurement circuits are constructed by using an ancilla circuit connecting two data qubits, which can be realised even in devices with minimal connectivity, such as the heavy-hexagonal layout [57, 56].
Figure 4: Depth-2 parity check measurement circuits. For the circuits to be bias-preserving, the CNOT gates in the circuits need to be also bias-preserving. Shown are the circuits for (a) $Z_{1}Z_{2}$, (b) $X_{1}X_{2}$, and (c) $Z_{1}X_{2}$ checks circuits, where the ancilla qubit is reset and later measured in the $Z$, $X$ and $Y$ bases, respectively.
@@
For a circuit to be bias preserving, it can be constructed using only gates that do not change the error type under conjugation. For example, one way to construct bias-preserving circuits is to use only CNOT gates in the measurement circuits as they only propagate errors to others with the same Pauli type. For the X3Z3 code, we need to construct bias-preserving circuits for three different kinds of parity checks: $ZZ$, $XX$ and $ZX$. All these three circuits can be constructed using only CNOT gates as shown in Fig. 4. As depicted, these circuits also include resets and measurements of the ancilla qubit in the three Pauli bases, i.e., $Z$, $X$ and $Y$ bases for $ZZ$, $XX$ and $ZX$ checks, respectively. Note that the circuits are still phase-bias preserving even if the resets/measurements in the $X$ and $Y$ bases are compiled in terms of resets/measurements in the $Z$ basis with additional single-qubit Clifford gates. This is because a $Z$ error on any qubit at any time step in the compiled circuits still propagates to data qubits as a $Z$ error, or as a check operator about to be measured.
@@
The parity-check circuits shown in Fig. 4, however, may not preserve the noise bias if errors happen during the CNOT gates. This is because a $Z$ error occurring on the target qubit during the application of a conventional (non-bias preserving) CNOT gate may result in a combination of $Z$ and $X$ errors on the target qubit after the CNOT gate [58]. One way to remedy this problem is to use bias-preserving CNOT gates [58]. These gates, however, are available only in specific platforms such as cat qubits. Here, we propose a more general approach to constructing $Z$-bias preserving parity check circuits without requiring the bias-preserving gates. While our proposed circuits can preserve only a specific type of noise bias, i.e., the $Z$ bias, these circuits are built using conventional (which may not be phase-bias preserving) two-qubit gates and thus can be implemented in many different architectures.
@@
Figure 5: Parity-check measurement circuits built from conventional two-qubit gates and constructed in ways that the $Z$-noise bias on the data qubits is preserved under mid-gate errors. Shown are (a) $ZZ$, (b) $XX$, and (c) $Z_{1}X_{2}$ check circuits. The circuits in (a) and (c) preserve the $Z$ noise bias on data qubits even in the presence of mid-gate errors. The circuit in (b) includes a classically controlled $X$ gate targeting one of the data qubits, which is non-$Z$ bias preserving. Data qubits are labelled by Data1 and Data2, while the middle qubit in each circuit is a measurement ancilla. In (c), the first equality can be checked by noting that the depth-3 circuit before and after the mid-circuit measurement performs a next-nearest-neighbour CZ gate, if the measurement ancilla is initially in the $\ket{+}$ state. The second equality can be checked by commuting all gates past the measurement. The grey measurements can be optionally included to provide flag information.
Our proposed phase-bias preserving parity-check circuits are depicted in Fig. 5. For the $ZZ$ and $ZX$ check measurement circuits, they are constructed using either CZ gates or CNOT gates targeting ancilla qubits, which propagate Pauli errors on the ancilla qubits only as $Z$ errors on the data qubits. In other words, these circuits do not leave any residual Pauli $X$ or $Y$ errors occurring with probability $\mathcal{O}(p_{Z})$ and therefore do not degrade the $Z$ noise bias on the data qubits. This is true even if we consider mid-gate Pauli errors (see Supplementary Sec. II in the Supplementary Information [61] for details). The $XX$ measurement circuit shown in Fig. 5(b), however, is not fully $Z$-bias preserving in the presence of mid-gate errors. This circuit degrades the $Z$ bias on one of the data qubits, i.e., the data qubit that is the target of the classically-controlled $X$ gate. (Note that instead of implementing the classically controlled $X$ gate, we may alternatively perform a CNOT gate, with the control being the ancilla qubit and target being data qubit 1, immediately preceding the ancilla measurement.) However, since the $XX$ parity-check circuit shown in Fig. 5(b) degrades the $Z$-bias on only one data qubit, it has a better $Z$-bias-preserving performance compared to the $XX$ check circuit in Fig. 4(b), which degrades the $Z$-bias on both data qubits due to the fact that both data qubits are used as target qubits of the CNOT gates in the circuit.
We now elaborate on how the parity-check circuits work. We begin by noting that while the $ZZ$ parity check measurements [Fig. 5(a)] are performed by reading out the ancilla qubits like in standard syndrome extraction circuits, the $XX$ and $ZX$ parity values [Figs. 5(b) and 5(c)] are obtained by measuring the data qubits after some two-qubit gates are applied on them. These two-qubit gates are required for the parity measurements so as not to reveal each individual data qubit’s state. For the $XX$ parity-check circuit, the two mid-circuit $X$-measurements in Fig. 5(b), are equivalent to reading out the $X_{1}X_{a}$ and $X_{2}X_{a}$ parity values, where the Pauli operator $X_{a}$ acts on the ancilla qubit which is prepared in the |$\ket{0}$ state. Hence, the product of these two measurement outcomes is the required $X_{1}X_{2}$ parity check outcome. We can disentangle the ancilla qubit by performing a final $Z_{a}$ measurement (which requires correcting a potential bit flip on one of the data qubits with the classically controlled $X$ gate), or by using a CNOT gate (not shown) before the $Z_{a}$ measurement (see Supplementary Sec. II of the Supplementary Information [61] for details).
The mixed-type $Z_{1}X_{2}$ parity circuit can be constructed using CZ gates between the data qubits, as shown in Fig. 5(c). If the connections between the data qubits are not available in the device such as in hardware with a heavy-hexagonal lattice, these CZ gates have to be implemented using ancilla qubits in-between the data qubits. We show such an implementation of next-nearest-neighbour CZ gates in the first equality of Fig. 5(c) (this is adapted from Ref. [56]). This circuit also includes optional measurements, coloured in grey, which provide “flag information” for detecting $Z$ errors on the ancilla qubit that may have propagated to the data qubits. Such flag measurements can assist in decoding [57]. In the second equality of Fig. 5(c), we reset the ancilla qubit in the |$\ket{0}$ state instead of the $\ket{+}$ state, allowing for a shorter-depth circuit to implement the $Z_{1}X_{2}$ measurement. This comes at the cost of removing one flag measurement. To see that this circuit works as intended, we can commute the two-qubit gates past the $X_{2}$ measurement, resulting in a $Z_{1}Z_{a}X_{2}$ parity measurement. From the result of this measurement, together with the fact that $Z_{a}$ value is set to be $+1$ at the ancilla reset, we can then infer the value of $Z_{1}X_{2}$.
We emphasise that, even if the $X$ measurements in the parity-check circuits above are implemented using $Z$ measurements sandwiched by Hadamard gates, this does not degrade the bias, since a $Z$ error occurring between those Hadamard gates is harmless. Specifically, a $Z$ error happening before the measurement is immediately absorbed by the $Z$ measurement without flipping its outcome, and a $Z$ error after the measurement is a stabiliser of the state (up to a sign), so does nothing. Moreover, the circuit will not change the noise bias significantly even if the Hadamard gates are noisy, since in many architectures, the single-qubit gate errors are not the predominant error source and are usually much smaller than the two-qubit gate errors [65, 66, 67, 68].
In Supplementary Sec. II of the Supplementary Information [61], we show in more detail how the parity check circuits preserve the noise bias. We also show that the phase-bias preserving parity-check circuits presented in this paper have optimal circuit depth.
II.7 Memory experiment simulation